# GraphPFN: A Prior-Data Fitted Graph Foundation Model

### v1

@@ 4 Graph prior
As discussed above, GraphPFN is based on the prior-data fitted networks (PFNs) framework (Müller et al., 2021). In this approach, the model is pretrained on a large number of synthetic graph datasets sampled from a chosen prior distribution. Because the pretrained model aims to approximate the posterior predictive distribution, it is crucial to design a high-quality and realistic prior. In the following sections, we describe the prior used for pretraining GraphPFN. First, we explain our method for generating realistic graph structures. Then, we describe how we use these graphs to generate node attributes and targets.
@@ 4.1 Structure generation
Our main aim is to generate graph structures similar to real-world graphs. After examining graphs from a range of graph machine learning datasets, we find that most of them exhibit strong cluster (community) structure, which in general is a common feature of real-world graphs (Girvan & Newman, 2002). Thus, as the basis for our graph generation process, we use the degree-corrected stochastic block model (SBM) (Karrer & Newman, 2011) that can generate graphs with community structure. However, we find that graphs generated from the degree-corrected SBM exhibit too “clean”-looking and well-defined clusters that can be visualized as several well-separated “balls”, while clusters in real-world graphs are often much more “rough”-looking, with more complex shapes and often overlapping with each other. Thus, to obtain graph structures similar to real-world ones, we design a novel method that combines multiple SBMs. First, we generate several first-level graphs from SBMs with different parameters. Then, we generate a second-level graph from another SBM, such that this second-level graph has the number of nodes equal to the sum of the numbers of nodes in all first-level graphs. We then randomly assign each node from the first-level graphs to a unique node in the second-level graph, thus essentially constructing a bijection $f$ between first-level and second-level graph nodes. Then, we transfer each edge from the first-level graphs to the second-level graph by creating a new edge in the second-level graph between the corresponding nodes, i.e., if there was an edge between nodes $u$ and $v$ in the first-level graphs, then we create an edge between the nodes $f(u)$ and $f(v)$ in the second-level graph. The obtained second-level graph with additional edges combines multiple graphs generated from different SBMs and exhibits clusters of nodes with complex shapes and overlaps that we were looking for.
@@
Further, we observe that most graphs from graph benchmarks exhibit core-periphery structure, where the core is composed of multiple relatively dense clusters, but there are also many low-degree peripheral nodes. Such structure is known to be common for real-world networks (Zhang et al., 2015). While our method of combining multiple graphs generated from SBM produces realistic node clusters, it produces a relatively small number of peripheral nodes. Thus, we augment the approach discussed above with a preferential attachment process (Price, 1965; 1976). Specifically, we run a modified Barabási–Albert (BA) process to add additional low-degree nodes to the graph (i.e., we use the graph obtained thus far as the initialization for the BA process). While the original BA process (Albert & Barabási, 2002) fixes the degree of newly-created nodes for the entire process, we treat this degree as a random variable and generate new nodes with different initial degrees.
@@
Our method has many hyperparameters such as the number and size of blocks for SBMs or their degree sequences. Similar to prior works on PFNs (Hollmann et al., 2023; 2025; Qu et al., 2025), we define probability distributions for each hyperparameter and sample new hyperparameters for each synthetic graph, which allows us to generate a diverse graphs. At the same time, we can easily set bounds for sizes, densities, or maximum degrees of the generated graphs, allowing us to ensure that the graphs fit the desirable constraints.
Since we need many thousands of synthetic graphs for training, the efficiency of generation becomes a concern. We utilize the implementations of the degree-corrected SBM and the BA process from the graph-tool library (Peixoto, 2017), which are highly efficient and allow even a single CPU core to generate multiple graphs in a second according to our process.
We provide example visualizations of several graphs generated by our process in Appendix C.
Table 1: The key statistics of the considered graph datasets.
tolokers-2 11,75811{,}758 519,000519{,}000 $16$ $88.3$ cls. 2 no tabular
artnet-exp 50,40550{,}405 280,348280{,}348 $75$ $11.1$ cls. 2 no tabular
hm-prices 46,56346{,}563 10,730,99510{,}730{,}995 $41$ $460.9$ reg. N/A no tabular
city-roads-M 57,07357{,}073 107,104107{,}104 $26$ $3.8$ reg. N/A yes tabular
artnet-views 50,40550{,}405 280,348280{,}348 $50$ $11.1$ reg. N/A no tabular
avazu-ctr 76,26976{,}269 10,984,07710{,}984{,}077 $260$ $288.0$ reg. N/A no tabular
twitch-views 168,114168{,}114 6,797,5576{,}797{,}557 44 $80.9$ reg. N/A no tabular
questions 48,92148{,}921 153,540153{,}540 $301$ $6.3$ cls. 2 no text-based
wiki-cs 11,70111{,}701 215,603215{,}603 $300$ $36.9$ cls. 10 yes text-based
pubmed 19,71719{,}717 44,32444{,}324 $500$ $4.5$ cls. 33 yes text-based
@@ 4.2 Attribute generation
We generate features and targets for synthetic graphs with a neural structural causal model (SCM) that extends the MLP-based SCM of Qu et al. (2025); Hollmann et al. (2023). As a starting point, we follow the TabICL protocol: we sample an MLP architecture (number of layers, dimension of hidden layers, activation function) and its weights at random, draw random inputs, propagate them through the network, and then designate a random subset of neurons as observed features and another random neuron as target, leaving the rest as latent variables. This yields a broad family of causal mechanisms in which features and targets can depend on each other and on latent confounders. We refer to Qu et al. (2025); Hollmann et al. (2023) for further details.
To make attributes also depend on the graph structure, we extend this SCM in two complementary ways. First, we introduce a mixture of MLP and GNN neurons. For each dataset, we sample a mixing probability $p\in\{0.0,0.1,\ldots,0.9,1.0\}$. At every hidden layer, we compute both an MLP transformation and a GNN layer. Each neuron is then independently assigned to be MLP-type or GNN-type, with probabilities 1p1-p and $p$, respectively, and its value is taken from the corresponding transformation. This mechanism controls how strongly the generated variables depend on the graph. Second, with a given probability, we optionally augment the random inputs with Laplacian positional encodings (LapPE) (Dwivedi et al., 2020; Belkin & Niyogi, 2001) to further integrate graph structure into the data generation process.
Together, the mixed MLP/GNN neurons and optional LapPE inject graph information into the SCM while remaining close to the tabular prior. When $p=0$ and LapPE is not used, the procedure reduces to the TabICL-style tabular SCM, while larger $p$ and the inclusion of LapPE increase the influence of graph structure on both features and targets.
@@ 5.1 Experimental setup
In terms of dataset selection, we closely follow the experimental setup of G2T-FM (Eremeev et al., 2025). We evaluate two collections of datasets: (i) real-world datasets from the recently proposed GraphLand benchmark (Bazhenov et al., 2025); and (ii) some of the classic graph datasets. Together, these datasets cover node classification and regression, come from diverse application domains, include both homophilous and heterophilous graphs, span a range of densities and other graph structural properties. Table 1 lists datasets used in our study and summarizes their statistics. For all datasets, we use 10%/10%/80% train/validation/test splits. Due to current limitations of TFMs, we restrict our study to small- and medium-scale datasets and exclude classification tasks with more than 10 classes. [^3] In principle, TFMs can handle more than 10 classes via schemes such as error-correcting output codes, as proposed in Hollmann et al. (2025). For simplicity, we focus on datasets with at most 10 classes that are natively supported by existing TFMs.
@@ Methods
  • Classic GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GT (Shi et al., 2021). Following Platonov et al. (2023), we augment these models with residual connections (He et al., 2016), layer normalization (Ba et al., 2016), and MLP blocks, which have been shown to substantially improve the performance of classic GNNs (Luo et al., 2024; 2025). We perform extensive hyperparameter tuning for these models.
@@
  • G2T-FM (Eremeev et al., 2025) with TabPFNv2 (Hollmann et al., 2025) and LimiX (Zhang et al., 2025) as the backbones. To the best of our knowledge, these are the strongest publicly available graph foundation models for node-level tasks with non-textual features.
We evaluate graph foundation models in both in-context learning (ICL) and full finetuning (FT) settings. We also optionally augment GraphPFN with Laplacian positional encodings (LapPE) (which are always used by G2T-FM). For further details, please refer to Appendix D.
Table 2: Results on GraphLand datasets under the RL data split.
@@
tolokers-2 artnet-exp hm-prices city-roads-M artnet-views city-reviews avazu-ctr twitch-views
LightGBM-NFA $56.34\pm 0.06$ $46.13\pm 0.03$ $70.84\pm 0.04$ $61.18\pm 0.03$ $56.10\pm 0.02$ $78.53\pm 0.01$ $31.71\pm 0.01$ $60.14\pm 0.01$
GCN $56.27\pm 0.29$ $44.86\pm 0.34$ $68.02\pm 0.40$ $58.82\pm 0.24$ $56.03\pm 0.24$ $77.81\pm 0.14$ $32.00\pm 0.15$ $75.51\pm 0.05$
GraphSAGE $54.43\pm 0.32$ $45.14\pm 0.34$ $70.00\pm 0.70$ $59.44\pm 0.26$ $49.32\pm 0.86$ $78.17\pm 0.09$ $31.44\pm 0.15$ $66.29\pm 0.31$
GAT $57.41\pm 0.80$ $45.06\pm 0.49$ $72.07\pm 1.16$ $59.86\pm 0.19$ $53.60\pm 0.23$ $77.74\pm 0.20$ $32.63\pm 0.16$ $72.89\pm 0.25$
GT $56.98\pm 0.53$ $46.41\pm 0.68$ $69.44\pm 0.89$ $59.55\pm 0.27$ $53.37\pm 0.43$ $77.34\pm 0.20$ $31.11\pm 0.47$ $72.13\pm 0.13$
G2T-TabPFNv2 (ICL) $60.42\pm 0.27$ $45.84\pm 0.03$ $66.68\pm 0.09$ $60.47\pm 0.04$ $58.75\pm 0.15$ $77.46\pm 0.10$ $26.38\pm 0.07$ $70.00\pm 0.06$
G2T-LimiX (ICL) $61.48\pm 0.30$ $48.43\pm 0.18$ $74.96\pm 0.06$ $64.53\pm 0.07$ $60.95\pm 0.10$ $77.72\pm 0.54$ $32.39\pm 0.14$ $71.08\pm 0.07$
G2T-TabPFNv2 (FT) $57.65\pm 1.92$ $47.31\pm 0.59$ $71.05\pm 0.91$ $63.08\pm 0.28$ $60.29\pm 0.13$ $79.12\pm 0.21$ $28.52\pm 0.43$ $74.06\pm 0.16$
G2T-LimiX (FT) $61.17\pm 0.49$ $49.88\pm 0.13$ $76.32\pm 0.17$ $65.87\pm 0.10$ $62.12\pm 0.10$ $80.13\pm 0.05$ $33.94\pm 0.34$ $73.16\pm 0.40$
GraphPFN (ICL) $58.95\pm 0.02$ $49.97\pm 0.01$ $68.02\pm 0.02$ $58.93\pm 0.02$ $58.87\pm 0.01$ $79.52\pm 0.01$ $21.67\pm 0.05$ $59.95\pm 0.08$
GraphPFN + LapPE (ICL) $61.18\pm 0.11$ $49.68\pm 0.04$ $68.56\pm 0.10$ $58.61\pm 0.41$ $60.31\pm 0.09$ $79.73\pm 0.08$ $23.33\pm 0.47$ $63.66\pm 0.77$
GraphPFN (FT) $60.77\pm 0.72$ $53.08\pm 0.23$ $79.15\pm 0.33$ $65.65\pm 0.20$ $64.15\pm 0.21$ OOM OOM OOM
GraphPFN + LapPE (FT) $61.83\pm 0.28$ $53.38\pm 0.18$ $78.58\pm 0.37$ $64.09\pm 0.38$ $64.39\pm 0.13$ OOM OOM OOM
Table 3: Results on other datasets.
facebook amazon-ratings questions wiki-cs pubmed
GCN $91.26\pm 0.19$ $41.43\pm 0.46$ $15.42\pm 0.63$ $81.74\pm 0.20$ $85.46\pm 0.18$
GraphSAGE $91.12\pm 0.21$ $40.07\pm 0.50$ $16.55\pm 0.61$ $81.50\pm 0.26$ $86.04\pm 0.26$
GAT $92.61\pm 0.20$ $40.67\pm 0.53$ $16.75\pm 0.63$ $82.25\pm 0.26$ $84.81\pm 0.22$
GT $91.71\pm 0.21$ $41.56\pm 0.38$ $14.03\pm 0.86$ $82.54\pm 0.20$ $84.95\pm 0.18$
G2T-TabPFNv2 (ICL) $90.56\pm 0.12$ $40.63\pm 0.19$ $16.49\pm 0.16$ $76.61\pm 0.57$ $88.80\pm 0.25$
G2T-LimiX (ICL) $91.29\pm 0.14$ $44.10\pm 0.16$ $15.31\pm 0.77$ $79.99\pm 0.28$ $88.96\pm 0.18$
G2T-TabPFNv2 (FT) $91.73\pm 0.28$ $44.71\pm 0.32$ $19.07\pm 0.53$ $79.70\pm 0.31$ $90.46\pm 0.11$
G2T-LimiX (FT) $92.16\pm 0.18$ $45.67\pm 0.35$ $20.19\pm 0.30$ $82.24\pm 0.31$ $89.91\pm 0.48$
GraphPFN (ICL) $90.23\pm 0.03$ $43.71\pm 0.03$ $12.30\pm 0.05$ $77.29\pm 0.03$ $89.76\pm 0.02$
GraphPFN + LapPE (ICL) $91.45\pm 0.16$ $42.64\pm 0.22$ $9.62\pm 0.16$ $79.13\pm 0.37$ $89.41\pm 0.09$
GraphPFN (FT) $93.06\pm 0.09$ $46.18\pm 0.17$ $20.50\pm 0.83$ $82.17\pm 0.22$ OOM
GraphPFN + LapPE (FT) $92.97\pm 0.19$ $45.78\pm 0.41$ $18.20\pm 3.51$ $82.15\pm 0.33$ OOM

### v2

@@ 4 Graph prior
As discussed above, GraphPFN is based on the prior-data fitted networks (PFNs) framework (Müller et al., 2022). In this approach, the model is pretrained on a large number of synthetic graph datasets sampled from a chosen prior distribution. Since the pretrained model aims to approximate the posterior predictive distribution, it is crucial to design a diverse, high-quality, and realistic prior. In this section, we describe the prior used for pretraining GraphPFN. First, we explain our method for generating realistic graph structures. Then, we describe how we use these graphs to generate node attributes and targets.
@@ 4.1 Structure generation
Our main aim is to generate graph structures similar to real-world graphs. After examining graphs from a range of graph machine learning datasets, we find that most of them exhibit strong clustering (community) structure, which in general is a common feature of real-world graphs (Girvan & Newman, 2002). Thus, as the basis for our graph generation process, we use the degree-corrected stochastic block model (SBM) (Karrer & Newman, 2011), which can generate graphs with community structure. However, we find that graphs generated from the degree-corrected SBM exhibit clusters that appear overly regular and well-defined, resembling separated spheres. In contrast, clusters in real-world graphs are often more irregular, with more complex shapes and frequent overlaps. Thus, to obtain graph structures similar to real-world ones, we design a novel method that combines multiple SBMs. First, we generate several first-level graphs from SBMs with different parameters. Then, we generate a second-level graph from another SBM, such that this second-level graph has the number of nodes equal to the total number of nodes across all first-level graphs. We then randomly assign each node from the first-level graphs to a unique node in the second-level graph, thus essentially constructing a bijection $f$ between first-level and second-level graph nodes. Then, we transfer each edge from the first-level graphs to the second-level graph by creating a new edge in the second-level graph between the corresponding nodes, i.e., if there was an edge between nodes $u$ and $v$ in the first-level graphs, then we create an edge between nodes $f(u)$ and $f(v)$ in the second-level graph. The obtained second-level graph with additional edges combines multiple graphs generated from different SBMs and exhibits clusters of nodes with complex shapes and overlaps that we aimed to capture.
@@
Further, we observe that most graphs from graph benchmarks exhibit a core-periphery structure, where the core is composed of multiple relatively dense clusters, but there are also many low-degree peripheral nodes. Such a structure is known to be common in real-world networks (Zhang et al., 2015). While our method of combining multiple graphs generated from SBMs produces realistic node clusters, it produces a relatively small number of peripheral nodes. Thus, we augment the approach discussed above with a preferential attachment (PA) process (Price, 1965; 1976; Albert & Barabási, 2002). Specifically, we use the graph obtained thus far as the initialization for the PA process. We sequentially add low-degree nodes and connect them to the previous ones with probabilities proportional to their degrees. The initial degree of each new node is chosen randomly.
@@
Our method has many hyperparameters such as the number and size of blocks for SBMs or their degree sequences. Similar to prior works on PFNs (Hollmann et al., 2023; 2025; Qu et al., 2025; 2026), we define probability distributions for each hyperparameter and sample new hyperparameters for each synthetic graph, which allows us to generate diverse graphs. At the same time, we can easily set bounds for sizes, densities, or maximum degrees of the generated graphs, allowing us to ensure that the graphs fit the desired constraints.
We provide example visualizations of several graphs generated by our process in Appendix B.1. We also analyze how training on simpler random graphs affects the downstream performance (Appendix B.3). Finally, we inspect the alignment of various characteristics between real and synthetic datasets in Appendix B.2.
@@ 4.2 Attribute generation
We generate features and targets for synthetic graphs with a neural structural causal model (SCM) that extends the MLP-based SCM of Qu et al. (2025); Hollmann et al. (2023). As a starting point, we follow the TabICL protocol: we sample an MLP architecture (number of layers, dimension of hidden layers, activation function) and its weights at random, draw random inputs, propagate them through the network, and then designate a random subset of neurons as observed features and another random neuron as the target, leaving the rest as latent variables. This yields a broad family of causal mechanisms in which features and targets can depend on each other and on latent confounders. We refer to Qu et al. (2025); Hollmann et al. (2023) for further details.
To make attributes also depend on the graph structure, we extend this SCM in two complementary ways. First, we introduce a mixture of MLP and GNN neurons. For each dataset, we sample a mixing probability $p\in\{0.0,0.1,\ldots,0.9,1.0\}$ and a GNN type from $\{\mathrm{GCN},\mathrm{SAGE_{mean}},\mathrm{SAGE_{max}},\mathrm{SAGE_{min}},\mathrm{GT}\}$. At every hidden layer, we compute the outputs of both an MLP transformation and a GNN layer. Each neuron is then independently assigned to be MLP-type or GNN-type, with probabilities 1p1-p and $p$, respectively, and its value is taken from the corresponding transformation. This mechanism controls how strongly the generated variables depend on the graph. Second, with a $0.5$ probability, we augment the random inputs with Laplacian positional encodings (LapPE) (Dwivedi et al., 2020; Belkin & Niyogi, 2001) of dimension uniformly sampled from $\{1,\ldots,32\}$ to further integrate graph structure into the data generation process.
Together, the mixed MLP/GNN neurons and optional LapPE inject graph information into the SCM while remaining close to the tabular prior. When $p=0$ and LapPE is not used, the procedure reduces to the TabICL-style tabular SCM, while a larger value of $p$ and the inclusion of LapPE increase the influence of graph structure on both features and targets.
@@ 5.1 Experimental setup
In terms of dataset selection, we follow the experimental setup of G2T-FM (Eremeev et al., 2025). We evaluate two collections of datasets: (i) real-world datasets from the recently proposed GraphLand benchmark (Bazhenov et al., 2025); and (ii) some of the classic graph datasets. Together, these datasets cover node classification and regression, come from diverse application domains, include both homophilous and non-homophilous graphs, [^2] A graph is called homophilous if its edges tend to connect nodes with similar labels, see Newman (2003); Platonov et al. (2023a); Mironov & Prokhorenkova (2024) for details. and span a range of densities and other graph structural properties. Table 7 lists the datasets used in our study and summarizes their statistics. For all datasets, we use 10%/10%/80% train/validation/test splits. Due to current limitations of TFMs, we restrict our study to small- and medium-scale datasets and exclude classification tasks with more than 10 classes. [^3] In principle, TFMs can handle more than 10 classes via schemes such as error-correcting output codes, as proposed in Hollmann et al. (2025). For simplicity, we focus on datasets with at most 10 classes that are natively supported by existing TFMs.
@@ Methods
  • Classic GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). Following Platonov et al. (2023b), we augment these models with residual connections (He et al., 2016), layer normalization (Ba et al., 2016), and MLP blocks, which have been shown to substantially improve the performance of classic GNNs (Luo et al., 2024; 2025). We perform extensive hyperparameter tuning for these models.
  • Prior GFMs: OpenGraph (Xia et al., 2024), AnyGraph (Xia & Huang, 2024), GCOPE (Zhao et al., 2024), and TS-GNN (Finkelshtein et al., 2025), which are not based on the PFN framework.
@@
  • Recent PFN-based GFMs: G2T-FM (Eremeev et al., 2025) and TAG (Hayler et al., 2025). To the best of our knowledge, these are the strongest publicly available graph foundation models for node-level tasks with arbitrary features.
We evaluate GraphPFN in both in-context learning (ICL) and full finetuning (FT) settings. Other models are evaluated in ICL, FT, or both settings based on their official implementations. For G2T-FM and GraphPFN in the ICL setting, we additionally report ensembling (E) results obtained by averaging the predictions over 10 forward passes with different random seeds. Since TAG inherently employs ensembling, we also mark it with “(ICL, E)”, despite the exact ensembling implementation being different from G2T-FM and GraphPFN. For the FT setting, we finetune the entire model, following Eremeev et al. (2025); Rubachev et al. (2025). Inspired by Abboud et al. (2020); Sato et al. (2021); Kanatsoulis et al. (2025), we additionally augment the input of GraphPFN with 8 random features during each forward pass in all experiments in this section. For further details on the experimental setup, please refer to Appendix D.
Table 1: Evaluation results on the GraphLand datasets under the RL (Random Low) data split. ICL stands for in-context learning, FT stands for fine-tuning, and E stands for ensembling. We report average precision for binary classification tasks and $R^{2}$ for regression tasks. We also report average rank over classification datasets in the “AR (cls)” column and over all datasets in the “AR (all)” column. During the computation of “AR (all)” we exclude all methods that do not support regression tasks, so that ranks are computed only over tables with no missing entries. For each column, we highlight first, second, and third best results with a color.
@@
artnet-exp city-reviews tolokers-2 artnet-views avazu-ctr city-roads-M hm-prices twitch-views AR (cls) AR (all)
LightGBM + NFA $46.13\pm 0.04$ $78.53\pm 0.01$ $56.34\pm 0.06$ $56.10\pm 0.02$ $31.71\pm 0.01$ $61.18\pm 0.03$ $70.84\pm 0.04$ $60.14\pm 0.01$ $\mathbf{11.0}$ $\mathbf{9.75}$
GCN $44.86\pm 0.36$ $77.81\pm 0.15$ $56.27\pm 0.31$ $56.03\pm 0.25$ $32.00\pm 0.16$ $58.82\pm 0.25$ $68.02\pm 0.42$ $75.51\pm 0.05$ $\mathbf{13.33}$ $\mathbf{10.25}$
GraphSAGE $45.14\pm 0.36$ $78.17\pm 0.10$ $54.43\pm 0.34$ $49.32\pm 0.91$ $31.44\pm 0.16$ $59.44\pm 0.27$ $70.00\pm 0.74$ $66.29\pm 0.32$ $\mathbf{12.33}$ $\mathbf{11.25}$
GAT $45.06\pm 0.52$ $77.74\pm 0.21$ $57.41\pm 0.85$ $53.60\pm 0.24$ $32.63\pm 0.17$ $59.86\pm 0.20$ $72.07\pm 1.22$ $72.89\pm 0.27$ $\mathbf{12.67}$ $\mathbf{9.5}$
GT $46.41\pm 0.71$ $77.34\pm 0.21$ $56.98\pm 0.55$ $53.37\pm 0.46$ $31.11\pm 0.49$ $59.55\pm 0.28$ $69.44\pm 0.94$ $72.13\pm 0.13$ $\mathbf{12.67}$ $\mathbf{10.88}$
OpenGraph (ICL) $15.16\pm 0.83$ $59.09\pm 0.72$ $40.38\pm 1.13$ $\mathbf{17.0}$
AnyGraph (ICL) $12.84\pm 0.93$ $63.71\pm 1.45$ $28.75\pm 3.56$ $\mathbf{18.33}$
TS-GNN (ICL) $20.44\pm 1.05$ $43.46\pm 5.17$ $38.54\pm 0.94$ $\mathbf{17.33}$
G2T-LimiX (ICL) $48.44\pm 0.23$ $77.81\pm 0.49$ $61.45\pm 0.33$ $60.94\pm 0.11$ $32.39\pm 0.13$ $64.53\pm 0.09$ $74.96\pm 0.07$ $71.08\pm 0.07$ $\mathbf{7.33}$ $\mathbf{7.5}$
GraphPFN (ICL) $50.85\pm 0.31$ $79.89\pm 0.10$ $60.33\pm 0.60$ $61.77\pm 0.23$ $30.41\pm 0.18$ $63.55\pm 0.29$ $75.38\pm 0.26$ $73.27\pm 0.13$ $\mathbf{5.67}$ $\mathbf{7.12}$
G2T-LimiX (ICL, E) $48.52\pm 0.29$ $77.97\pm 0.53$ $61.58\pm 0.32$ $60.96\pm 0.10$ $32.41\pm 0.14$ $64.69\pm 0.05$ $74.97\pm 0.06$ $71.30\pm 0.07$ $\mathbf{6.33}$ $\mathbf{6.5}$
TAG-TabPFNv2 (ICL, E) $47.87\pm 0.39$ $77.38\pm 0.29$ $59.33\pm 1.00$ $\mathbf{11.33}$
TAG-LimiX (ICL, E) $50.19\pm 0.46$ $78.87\pm 0.13$ $59.46\pm 1.05$ $\mathbf{7.0}$
GraphPFN (ICL, E) $51.49\pm 0.10$ $80.22\pm 0.04$ $60.98\pm 0.25$ $62.07\pm 0.04$ $30.46\pm 0.14$ $64.29\pm 0.12$ $75.82\pm 0.06$ $73.81\pm 0.04$ $\mathbf{3.0}$ $\mathbf{5.25}$
GCOPE (FT) $14.92\pm 1.56$ $67.16\pm 0.98$ $28.81\pm 1.28$ $\mathbf{17.33}$
G2T-LimiX (FT) $49.89\pm 0.18$ $80.14\pm 0.06$ $61.28\pm 0.70$ $62.08\pm 0.12$ $34.02\pm 0.27$ $65.82\pm 0.18$ $76.27\pm 0.24$ $73.69\pm 0.35$ $\mathbf{5.0}$ $\mathbf{4.25}$
GraphPFN (FT) $50.64\pm 1.27$ $80.49\pm 0.16$ $60.90\pm 0.98$ $63.90\pm 0.12$ $34.34\pm 0.50$ $66.04\pm 0.41$ $79.26\pm 0.43$ $78.17\pm 0.13$ $\mathbf{4.0}$ $\mathbf{2.75}$
G2T-LimiX (FT, E) $49.96\pm 0.09$ $80.15\pm 0.07$ $60.71\pm 0.57$ $62.11\pm 0.06$ $33.86\pm 0.33$ $65.96\pm 0.11$ $76.30\pm 0.19$ $73.62\pm 0.49$ $\mathbf{5.67}$ $\mathbf{4.38}$
GraphPFN (FT, E) $51.97\pm 0.40$ $80.89\pm 0.05$ $60.92\pm 0.56$ $65.12\pm 0.11$ $35.05\pm 0.23$ $67.27\pm 0.14$ $80.13\pm 0.37$ $78.50\pm 0.36$ $\mathbf{2.33}$ $\mathbf{1.5}$